The dimension of a rectangular shape is 5.9 by 4.1mm s measured with a veneer calliper whose minimum scale is 0.05 mm. The percentage error in its surface area is approximately.
1
Expert's answer
2020-06-03T12:11:04-0400

Explanations & Calculations


  • Fractional error of a measurement is "\frac{\text{error}}{\text{measurement}}"


  • In measuring instruments the least measurement is considered as the maximum error.
  • The effect of errors in taking measurements of several quantities, is affected in a way of fractional errors when a product or ratio of those quantities are considered.
  • With respect to this situation, l = 5.9 mm, b =4.1 mm & Area = l "\small \times" b


  • If the possible maximum error is noted by "\delta" ,

"\qquad\qquad
\begin{aligned}
\small \frac{\delta(Area)}{Area} &= \small \frac{\delta(l)}{l} +\frac{\delta(b)}{b}
\end{aligned}"


  • Therefore, for this question,

"\qquad\qquad
\begin{aligned}
\small \frac{\delta(Area)}{Area} &= \small \frac{0.05mm}{5.9mm} +\frac{0.05mm}{4.1mm}\\
&= \small \frac{50}{2419}\\
& =\small 0.020669 
\end{aligned}"

  • Therefore, percentage error is 0.020669 "\times" 100% = 2.067%